The distance between the first and fifth minima of a single-slit diffraction pattern is 0.500 mm with the screen 39.0 cm away from the slit, when light of wavelength 550 nm is used. Find the slit width.
Answer #1

For minima:

y = mD\lambda/a : where a =slit width ,m= order of minima,\lambda= wavelength,D= slit-screen distance

And here y5 - y1 = 5*(D\lambda/a) - (D\lambda/a) = 4*(D\lambda/a) = 0.5 mm = 4*(39cm*550nm/a)

=> a = 4*0.39*2000*550nm = 1716000 nm = 1.716 mm

So, slit width = 1.716 mm

Please do rate with points, it'll help me